# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v3

@@ D.3 Consistency-Reactivity Inequalities
Proposition 1 (Consistency-Reactivity Inequalities). Let $\mathcal{L}$ be a non-linear convex function measuring the prediction error with respect to demonstrations. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Similarly, let $\mathcal{P}^{+}\coloneqq\{a_{t-h:t-1}\}$ and $\mathcal{P}^{-}\coloneqq\{a_{t-h-d:t-h-1}\}$. Let $C\coloneqq\mathcal{P}^{+}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{P}^{-}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
@@
$\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d}).$
@@
$=\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P(\hat{s}_{t-h-d+1:}^{t-h}|s_{t-h-d-c:}^{t-h-d},a_{t-h-d:}^{t-h-1})\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P_{f}^{d}\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
@@
$\hat{P}_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1}).$
@@
$P_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}s_{t-h-d+1:t-h}\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P({s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1}).$
@@
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P_{f}^{d}\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P_{f}^{d}\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\hat{P}_{\neq g_{t^{\prime}}}\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[P_{f}^{d}\mathcal{L}(\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
@@
$\left[P_{f}^{d}\left.P_{f}^{d}\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\left.P_{\neq g_{t^{\prime}}}\left.P_{f}^{d}\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+P_{f}^{d}\hat{P}_{\neq g_{t^{\prime}}}\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]$
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off, first, define $P_{{\hat{s}=s}}^{f}\coloneqq P({s}_{t-h-d+1:}^{t-h}|s_{t-h-d},a_{t-h-d:}^{t-h-1})$. Then,
@@
$\left[P_{f}^{d}\left.P_{f}^{d}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\left.P_{\neq g_{t^{\prime}}}\left.P_{\hat{s}=s}^{f}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+P_{f}^{d}\hat{P}_{\neq g_{t^{\prime}}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]$
$+P_{\neq g_{t^{\prime}}}P_{f}^{d}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\neq{s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]$
$+P_{\neq g_{t^{\prime}}}P(\hat{s}_{t-h-d+1:}^{t-h}\neq s_{t^{\prime}}|s_{t-h-d},a_{t-h-d:}^{t-h-1})$
@@
For the match terms, we use the fact that $P_{f}^{d}\leq 1$ and
$P_{{\hat{s}=s}}^{f}=P(\hat{s}_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\leq 1$. For the mismatch terms, we use the definition of $\epsilon_{f}$ and Lemma 5. Then, we continue:
@@
$+P_{\neq g_{t^{\prime}}}\left[P_{f}^{d}\epsilon_{f}+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}}\epsilon_{f}\right]+P_{f}^{d}\left[\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{f}\right].$ (Simulation does not match reality)
We simplify the mismatch terms further:
$\leq P_{\neq g_{t^{\prime}}}\left[P_{f}^{d}\epsilon_{f}+(1-P_{f}^{d})\epsilon_{f}\right]+P_{f}^{d}\left[\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{f}\right]$
$=P_{\neq g_{t^{\prime}}}\epsilon_{f}+P_{f}^{d}\hat{P}_{\neq g_{t^{\prime}}}\epsilon_{f}$
$=P_{\neq g_{t^{\prime}}}\epsilon_{f}+\left.P_{f}^{d}\left[(1-P_{f}^{d})\right]\epsilon_{f}\right.$
$=(1-P_{f}^{d})\epsilon_{f}+\left.P_{f}^{d}\left[(1-P_{f}^{d})\right]\epsilon_{f}\right.$
$=\epsilon_{f}\cdot\left[1-P_{f}^{d}+P_{f}^{d}-P_{f}^{2d}\right]$
$=\epsilon_{f}\cdot\left[1-P_{f}^{2d}\right].$
@@
$=\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P(g_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\pi_{(c+d,h)}+\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:}^{t-h-c-1},\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\pi_{(c+d,h)}^{t},\pi^{*})\Big{|}C\right].$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(P_{b}^{d}\pi_{(c+d,h)}+\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:}^{t-h-c-1},\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\pi_{(c+d,h)}^{t},\pi^{*})\Big{|}C\right].$
@@
$\hat{P}_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:t-h-c-1}\\ \text{not all }g_{t^{\prime}}\end{subarray}}P(\hat{s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c}).$
@@
$P_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}s_{t-h-d-c:t-h-c-1}\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P({s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c}).$
@@
$\leq\min_{\pi_{(c+d,h)}}\left.\mathbb{E}_{G}\left[\mathcal{L}(P_{b}^{d}\pi_{(c+d,h)}(a_{t}|s_{t-h-c:}^{t-h},g_{t-h-d-c:}^{t-h-c-1},a_{t-h:}^{t-1})+\right.\right.$
$\hskip 14.22636pt\left.\hat{P}_{\neq g_{t^{\prime}}}\pi_{(c+d,h)}(a_{t}|s_{t-h-c:}^{t-h},\hat{s}_{t-h-d-c:}^{t-h-c-1},a_{t-h:}^{t-1}),\pi^{*})\mid C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[P_{b}^{d}\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\mid C\right]+$
@@
$\left[P_{b}^{d}\left.P_{b}^{d}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right]\right.\right.$
$+P_{\neq g_{t^{\prime}}}\left.P_{b}^{d}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\neq s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+P_{b}^{d}\hat{P}_{\neq g_{t^{\prime}}}\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}=s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off and defining $P_{\hat{s}=s}^{b}\coloneqq P(\hat{s}_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})$:

### v4

@@ E.3 Consistency-Reactivity Inequalities
Proposition 1 (Consistency-Reactivity Inequalities) Let $\mathcal{L}$ be a non-linear and non-negative convex function measuring the prediction error with respect to demonstrations. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Let $C\coloneqq\{a_{t-h:t-1}\}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\{a_{t-h-d:t-h-1}\}\cup\mathcal{S}^{-}$. For notational ease, let $\tau_{f}=\{t-h-d+1:t-h\}$ and let $\tau_{b}=\{t-h-d-c:t-h-c-1\}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
@@
$\alpha_{f}-\epsilon_{b}(1-\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})(P_{b}(t^{\prime})+\delta_{b}(t^{\prime})))\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$
and
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq\epsilon_{f}(1-\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})(P_{f}(t^{\prime})+\delta_{f}(t^{\prime})))-\alpha_{b}.$
@@
$=\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\hat{P}_{\mathrm{train}}(g_{t-h-d+1:}^{t-h}|s_{t-h-d-c:}^{t-h-d},a_{t-h-d:}^{t-h-1})\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
@@
$\hat{P}_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1}).$
@@
$P_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}s_{t-h-d+1:t-h}\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P_{\mathrm{test}}({s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1}).$
@@
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\right.$
$\left.\sum_{\begin{subarray}{c}\hat{s}_{t-h-d+1:t-h}\\ \text{not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h})+\hat{P}_{\neq g_{t^{\prime}}}\right.\left.\pi_{(c+d,h)}(a_{t}|\cdots,\hat{s}_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\mathcal{L}(\pi_{(c+d,h)}(a_{t}|\cdots,g_{t-h-d+1:}^{t-h}),\pi^{*})|C\right]$
@@
$\left[\prod_{t^{\prime}\in\tau_{f}}P_{f}^{\mathrm{test}}(t^{\prime})\left.\hat{P}_{f}(t^{\prime})\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\left.P_{\neq g_{t^{\prime}}}\left.\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\prod_{t^{\prime}\in\tau_{f}}P_{f}^{\mathrm{test}}(t^{\prime})\hat{P}_{\neq g_{t^{\prime}}}\mathbb{E}_{G^{\prime}}\left[\mathcal{L}(\pi_{(c+d,h)}(...\hat{s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]$
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off, first, define $\hat{P}_{{\hat{s}=s}}^{f}$ to be the probability that the inferred states at timesteps in $\tau_{f}$ are the states that were visited in reality. Then,
@@
$\left[\prod_{t^{\prime}\in\tau_{f}}P_{f}^{\mathrm{test}}(t^{\prime})\left.\hat{P}_{f}(t^{\prime})\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\left.P_{\neq g_{t^{\prime}}}\left.\hat{P}_{\hat{s}=s}^{f}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]\right.\right.$
$+\prod_{t^{\prime}\in\tau_{f}}P_{f}^{\mathrm{test}}(t^{\prime})\hat{P}_{\neq g_{t^{\prime}}}\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\right]$
$+P_{\neq g_{t^{\prime}}}\prod_{t^{\prime}\in\tau_{f}}\hat{P}_{f}(t^{\prime})\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,\hat{s}_{t-h-d+1:}^{t-h}=g_{t-h-d+1:}^{t-h}\neq{s}_{t-h-d+1:}^{t-h}),\pi^{*})\Big{|}...,s_{t-h-d+1:}^{t-h}\neq g_{t-h-d+1:}^{t-h}\right]$
$+P_{\neq g_{t^{\prime}}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}\neq s_{t-h-d+1:}^{t-h},g_{t-h-d+1:}^{t-h}|s_{t-h-d},a_{t-h-d:}^{t-h-1})$
@@
For the match terms, we use the fact that $\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})\leq 1$ and
$\hat{P}_{{\hat{s}=s}}^{f}=\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d+1:}^{t-h}={s}_{t-h-d+1:}^{t-h}|{s}_{t-h-d},a_{t-h-d:}^{t-h-1})\leq 1$. For the mismatch terms, we use the definition of $\epsilon_{f}$ and Lemma 5. Then, we continue:
@@
$+\prod_{t^{\prime}\in\tau_{f}}P_{f}^{\mathrm{test}}(t^{\prime})\left[(1-\hat{P}_{f}(t^{\prime}))\epsilon_{f}\right]+\prod_{t^{\prime}\in\tau_{f}}(1-P_{f}^{\mathrm{test}}(t^{\prime}))\left[\hat{P}_{f}(t^{\prime})+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}}\right]\epsilon_{f}.$ (Simulation does not match reality)
Recall that we write $P_{f}^{\mathrm{test}}(t^{\prime})$ as simply $P_{f}(t^{\prime})$. We simplify the mismatch terms further:
$\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})\left[(1-\hat{P}_{f}(t^{\prime}))\epsilon_{f}\right]+\prod_{t^{\prime}\in\tau_{f}}(1-P_{f}(t^{\prime}))\left[\hat{P}_{f}(t^{\prime})+\hat{P}_{\neq g_{t^{\prime}},s_{t^{\prime}}}\right]\epsilon_{f}$
$\leq\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})\left[(1-\hat{P}_{f}(t^{\prime}))\epsilon_{f}\right]+\prod_{t^{\prime}\in\tau_{f}}(1-P_{f}(t^{\prime}))\left[\hat{P}_{f}(t^{\prime})+(1-\hat{P}_{f}(t^{\prime}))\right]\epsilon_{f}$
$=\epsilon_{f}\cdot\left[1-\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})\hat{P}_{f}(t^{\prime})\right].$
$=\epsilon_{f}\cdot\left[1-\prod_{t^{\prime}\in\tau_{f}}P_{f}(t^{\prime})(P_{f}(t^{\prime})+\delta_{f}(t^{\prime}))\right].$
@@
$=\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\mathcal{L}(\hat{P}_{\mathrm{train}}(g_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\pi_{(c+d,h)}+\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:}^{t-h-c-1},\\ \text{ not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\pi_{(c+d,h)},\pi^{*})\Big{|}C\right].$
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\prod_{t^{\prime}\in\tau_{b}}\hat{P}_{b}(t^{\prime})\mathcal{L}(\pi_{(c+d,h)},\pi^{*})+\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:}^{t-h-c-1},\\ \text{ not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})\mathcal{L}(\pi_{(c+d,h)},\pi^{*})\Big{|}C\right].$
@@
$\hat{P}_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}\hat{s}_{t-h-d-c:t-h-c-1}\\ \text{not all }g_{t^{\prime}}\end{subarray}}\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c}).$
@@
$P_{\neq g_{t^{\prime}}}:=\sum_{\begin{subarray}{c}s_{t-h-d-c:t-h-c-1}\\ \text{ not all }g_{t^{\prime}}\end{subarray}}P_{\mathrm{test}}({s}_{t-h-d-c:}^{t-h-c-1}|{s}_{t-h-c}).$
@@
$\leq\min_{\pi_{(c+d,h)}}\mathbb{E}_{G}\left[\prod_{t^{\prime}\in\tau_{b}}\hat{P}_{b}(t^{\prime})\mathcal{L}(\pi_{(c+d,h)}(a_{t}|...,g_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\mid C\right]+$
@@
$\left[\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})\hat{P}_{b}(t^{\prime})\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right.\right.\right]$
$+P_{\neq g_{t^{\prime}}}\left.\prod_{t^{\prime}\in\tau_{b}}\hat{P}_{b}(t^{\prime})\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\neq s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
$+\prod_{t^{\prime}\in\tau_{b}}P_{b}(t^{\prime})\hat{P}_{\neq g_{t^{\prime}}}\left.\mathbb{E}\left[\mathcal{L}(\pi_{(c+d,h)}(...,\hat{s}_{t-h-d-c:}^{t-h-c-1}\neq g_{t-h-d-c:}^{t-h-c-1}=s_{t-h-d-c:}^{t-h-c-1}),\pi^{*})\Big{|}...,s_{t-h-d-c:}^{t-h-c-1}=g_{t-h-d-c:}^{t-h-c-1}\right.\right]$
@@
Now, we group all the terms into two - one representing where the learner’s simulation matches the reality and one where it does not. Continuing from where we left off and defining $\hat{P}_{\hat{s}=s}^{b}\coloneqq\hat{P}_{\mathrm{train}}(\hat{s}_{t-h-d-c:}^{t-h-c-1}={s}_{t-h-d-c:}^{t-h-c-1}|s_{t-h-c})$: